OpenStudy (anonymous):

(x-5)^2+(y+4)^2=64 I know the center is (5,-4) and radius is 8 but how do I get that answer?

hero (hero):

It comes from the fact that if given a circle of the form: (x - h)^2 + (y - k)^2 = r^2, then (h,k) is the center and r is the radius.

OpenStudy (anonymous):

Thank you Hero, could you please go into a little more detail on how to work the problem?

hero (hero):

There's not really much to work in this case. It mostly involves observation, recognition and understanding the components definition of the equation of a circle. Given it's definition, you identify its components. The center (h,k) and radius, r are part of the general components. You take those components and identify them specifically for any given equation of a circle.

hero (hero):

If given the equation of a circle (x + 3)^2 + (y - 2)^2 = 16, what is the center and radius? The center is (-3, 2) and the radius is 4

OpenStudy (anonymous):

thank you so much! I get what you are saying! thank you!